OpenStudy (anonymous):

Solve the following system of equations. 2a – 3b + c = 10 2a – 2b – 2c = 2 a + 3b + 2c = –1

OpenStudy (anonymous):

do u know how to use matrix method?

OpenStudy (anonymous):

first one a=-c/2+3b/2+5

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

second one a=b+c+1

OpenStudy (anonymous):

third one b=-2c/3-1/3

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

lauren are u still there

OpenStudy (anonymous):

or u can use this way 2a – 3b + c = 10 R1 2a – 2b – 2c = 2 R2 a + 3b + 2c = –1 R3 keep R1 R2' = R1-R2 (illumination method) = -b + 3c = 8 R3' = R1-2R3 (illumination method) = -9b -3c = 12 now the system changed to 2a-3b+c = 10 R1 -b+3c = 8 R2' -9b-3c = 12 R3' one more time, now we keep R1 and R2' R3''= R2'+R3' = -10b = 20 finally the system changed to 2a-3b+c = 10 R1 -b+3c = 8 R2' -10b=20 R3'' now u can get b from R3'', then use the back-substitution to find a and c The answer are: b=-2, c=2 , a= 1

OpenStudy (anonymous):

if u know how to use matrix method, it would be easier than this way